Fermi National Laboratory


Accelerator Update for December 6, 2004 - December 8, 2004

Monday December 6
The day shift began with Antiproton Source (Pbar) stacking, with the Tevatron (TeV) and with Recycler conducting studies, and with MiniBooNE taking beam.

Operations put the TeV into shot setup at 9:47 AM. At 10:56 AM, Operations established a 14x0 proton only study-store in the TeV.

Operations put the TeV, Main Injector (MI), and Pbar into shot setup at 1 PM. Over the afternoon, six pilot shots of antiprotons were sent to MI, but all fell out before they got to acceleration due to a low-level RF problem. Investigation showed a cable problem with the phase detector in the tunnel. Experts decided to access the tunnel to make repairs.

Pbar resumed stacking at 5:08 PM.

Operators began preparing MI for access at 5:10 PM.

Operators began preparing sector F in the TeV at 6:07 PM for an access. Two MI experts will search the Switchyard (SY) 120 line for a possible obstruction. In addition, the Collider Detector at Fermilab (CDF) and D-Zero (D0) both accessed their collision halls.

CDF came out of access at 7:34 PM.

The experts in F sector reported at 7:52 PM that they found a magnet with a bus problem. Operations checked with an expert who said it would take around 2.5 hours to repair. Since the MI repair would be completed in less time, the F sector work was put off.

D0 came out of access at 9:19 PM.

Operations began the TeV turn on at 9:25 PM.

Operations put the TeV into a dry squeeze at 11:30 PM.

MI experts reported at 11:44 PM that they replaced parts on the phase detectors, but they think the same parts would just burn out again. They began a search to find the parts to make a better fix.

Tuesday December 7
The midnight shift began with MI experts searching for parts, which put stacking, beam to MiniBooNE, and the TeV octupole studies on hold.

MI experts reported at 12:34 AM that the phase detector was fixed.

At 1:48 AM, operators reported no beam in Linac. Operators accessed the I- Source and found no lights or rack power. An expert came in and at 2:25 AM reported that the source had died earlier, but it hadn't been noticed because there were no users until MI came back on. More experts came in and switched to the H- Source at 3:04 AM.

Experts re-established Linac beam at 3:26 AM.

Operations began sending beam to MiniBooNE at 3:55 AM.

Operators attempted to send beam to Pbar, but a power supply problem (HV102) stopped them. Operations contacted an expert and got the supply back on.

TeV helix studies began at 4:13 AM.

Pbar began stacking at 4:24 AM.

The TeV lost its beam when the sector DA lead tripped off at 5:48 AM. The TeV abort kicker (C:PAK2) also tripped. Operations reported that this same thing had happened last Sunday. At 6:48 AM, Operations called in an expert.

Operations began a TeV shot setup at 9:52 AM.

Operations established store 3821 at 3:25 PM. Its initial luminosity was 10.03E30.

A MI expert began a slipstacking study at 3:30 PM.

At 6:44 PM, store 3821 fell out of the TeV and quenched at sector E1. A TeV expert was here to help discover the cause.

A Recycler expert began a study at 7 PM.

Pbar resumed stacking at 8:58 PM, but at very poor efficiency. An expert was called in.

TeV experts reported at 10:22 PM that they think the quench was due to an A17 separator spark.

Operations halted Pbar stacking so an expert could work on a septum.

Operations put the TeV through a dry squeeze at 11:35 PM.

Wednesday December 8
The midnight shift began with Accumulator work, with the TeV and Recycler conducting studies, and with MiniBooNE taking beam.

The septum work was completed and Pbar resumed stacking at 12:24 AM.

Recycler studies ended at 1:02 AM.

Pbar experts began orbit studies and P2 tuning, all parasitic to stacking, at 5:30 AM.

Operations loaded a 36x0 study-store into the TeV at 7:13 AM.

Plans for the Wednesday
The plan for today is to go into shot setup later this morning. There will be an access into TeV sector F and A3 tomorrow morning.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 12/8/2004   email Fermilab